The specific volume of a gas is 1.5 m³/kg. What volume does 7.5 kg of the gas occupy? What is the density of the gas?
1
Expert's answer
2021-10-28T08:52:26-0400

"v=1.5 \;m^3/kg \\

ρ= \frac{1}{v} \\

ρ= \frac{1}{1.5} = 0.67 \;kg/m^3 \\

m=7.5 \;kg \\

V = \frac{m}{ρ} \\

V = \frac{7.5}{0.67} = 11.19 \;m^3"

Density = 0.67 kg/m3

Volume = 11.19 m3